OpenStudy (anonymous):

Write a simplified polynomial expression to represent the area of the square tile, shown below.

OpenStudy (anonymous):



OpenStudy (anonymous):

Area = lw = (x-3)(x-3) = (x-3)^2 = x^2 - 6x + 9

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes :)